# SocialTrack: Multi-Object Tracking in Complex Urban Traffic Scenes Inspired by Social Behavior

### v2

@@ IV-E Comparison with Other State-of-the-Art Methods
TABLE III: Performance test of VACKF module.
Speed Adaptive Model CKF MOTA $\uparrow$ IDF1 $\uparrow$ VACKF 62.5 74.1 62.7 74.4 63.2 74.7
IV-F Ablation Experiments
UAVDT: In this study, ByteTrack[26] was used as the baseline, and the effectiveness of each proposed module was validated through systematic ablation experiments. Table II shows the improvements in overall tracking performance due to four key modifications. First, to enhance the detector’s ability to perceive small targets, we designed a specialized detector for efficient small-target feature perception. The experimental results indicate that this module improved MOTA from 56.1 to 62.5, MOTP from 75.2 to 76.0, and IDF1 from 68.6 to 74.1, with significant progress in all three key metrics. This demonstrates that the improvement in detection quality plays a crucial role in the overall tracking performance. Additionally, IDSW decreased by 128, indicating that more accurate target detection also helps maintain identity consistency in the subsequent association process. Building on the detection results, we introduced the VACKF, which combines the advantages of the velocity-adaptive model and cubature Kalman filter for nonlinear state modeling and prediction. This module increased MOTA to 63.2, MOTP slightly to 76.1, and IDF1 to 74.7, while further decreasing IDSW to 304. This result demonstrates that the adaptive filtering mechanism significantly reduces trajectory estimation errors and identity switching frequency in scenes with frequent occlusions or complex motion, thereby enhancing the robustness and continuity of the tracking system. Next, we introduced the GMCS, which leverages the motion prior information from neighboring trajectories in the group to provide robust relative state references for low-quality trajectories, further improving the accuracy of their state estimation. This module increased MOTA to 64.2, IDF1 to 75.4, and reduced IDSW to 286, confirming the value of group behavior information in weak observation scenarios. Finally, when the GMCS fails, the STMP uses historical trajectory information to predict the future state of low-quality trajectories, effectively enhancing the temporal robustness of the model. This module further increased MOTA to 64.6, IDF1 to 76.1, and reduced IDSW to 276, indicating its ability to maintain target identity continuity in complex scenes.
To address the limitations of classical Kalman filter in predicting nonlinear motion, we introduced several enhancements to the filtering process: a velocity-adaptive model was employed to capture the nonlinear dynamics of moving objects and was integrated into the CKF. As shown in Table III, the inclusion of the velocity-adaptive model led to an increase in MOTA to 62.7 and IDF1 to 74.4, indicating the model’s ability to adapt to non-uniform motion and effectively mitigate estimation biases caused by conventional linear prediction assumptions. Building on this, the integration of the CKF resulted in the complete VACKF model, which further improved MOTA to 63.2 and IDF1 to 74.7, representing the highest scores in this group of experiments. These results demonstrate that CKF’s capability to model nonlinear state transitions complements the adaptive motion modeling, enhancing the filter’s ability to cope with complex and dynamic scenarios. Overall, the experimental results confirm that the design of VACKF significantly improves state estimation accuracy, reduces trajectory jitter and identity switches, and enhances the overall stability and robustness of the tracking system under challenging conditions.
TABLE IV: Performance Improvement with Different Components added to ByteTrack.
@@ IV-F Ablation Experiments
TABLE V: Performance Improvement with Different Components added to BoT-SORT.
@@
MOT17: To further validate the generalization and robustness of the proposed modules, we integrated SocialTrack (excluding the specialized detector) into two mainstream multi-object tracking frameworks—ByteTrack[26] and BoT-SORT[21]—and conducted systematic ablation experiments on the MOT17 dataset while keeping the detector consistent (YOLOX-X). The experimental results are shown in Table IV and Table V.
In the ByteTrack framework, after adding VACKF, HOTA improved from 68.0 to 68.1, MOTA increased from 76.6 to 76.7, IDF1 rose from 79.3 to 79.7, and both IDR and IDP showed slight increases. This indicates that even under conditions of limited detection accuracy, VACKF can enhance identity preservation and overall stability in the tracking process through its nonlinear modeling capabilities and adaptive state estimation mechanism. After introducing the GMCS, the overall performance improved further, with IDF1 reaching 80.2, HOTA rising to 68.4, and IDSW decreasing from 159 to 156. By effectively modeling the motion consistency among individuals in the local neighborhood, this module demonstrated stronger identity continuity in scenarios involving occlusion and crossing motion. With the addition of the STMP, IDF1 further increased to 80.4, and IDSW significantly dropped to 139, demonstrating that this module can accurately predict states in situations with low observation quality or temporary target loss, thereby effectively mitigating identity switching and enhancing the system’s temporal robustness.
@@
In the BoT-SORT framework, a similar trend was observed, further confirming the good generalizability of the proposed method. Compared to the original BoT-SORT, VACKF did not lead to performance improvements, which can be attributed to the camera motion compensation already addressing the position prediction shortcomings. After introducing the GMCS, IDF1 improved from 81.4 to 81.7, IDP increased to 87.0, and IDSW significantly decreased to 123, indicating that this strategy can effectively integrate motion prior information into BoT-SORT’s feature modeling mechanism, improving association accuracy in weak observation scenarios. Finally, with the addition of the STMP, the IDF1 of BoT-SORT increased to 82.2, HOTA rose to 69.4, and IDSW further decreased to 122, showing a consistent performance improvement trend similar to ByteTrack. This module demonstrated robust and transferable performance enhancement across different frameworks, further proving the proposed method’s good cross-framework adaptability and potential for broad application.
V Conclusion
@@
This study proposes a novel multi-object tracking algorithm, SocialTrack, designed to address the challenges of tracking small targets in complex environments. Through a series of ablation and comparison experiments, we have demonstrated the effectiveness of the proposed improvements in enhancing tracking accuracy, robustness, and stability. Compared to current mainstream methods, SocialTrack leads in several metrics, particularly in core performance indicators such as MOTA (64.6%) and IDF1 (76.1%), achieving improvements of 9.3% and 4.8%, respectively, over the existing best methods, setting a new state-of-the-art. Additionally, SocialTrack shows superior performance in handling complex scenarios, particularly under frequent occlusions and dense targets, effectively reducing identity switching and target loss, demonstrating stronger robustness. In conclusion, SocialTrack significantly improves the overall performance of multi-object tracking in complex environments by systematically integrating multiple innovative modules. Furthermore, all innovative modules possess strong plug-and-play capabilities and compatibility, allowing them to be incorporated into existing tracking algorithms to further enhance performance, demonstrating the great potential of this method for practical applications. Therefore, SocialTrack provides an important reference and insight for future multi-object tracking research and engineering applications.

### v3

@@ 4.6 Ablation Experiments
Baseline 56.1 75.2 68.6 447 442 196
+SOFEPNet 62.5 76.0 74.1 319 645 174
+SOFEPNet+VACKF 63.2 76.1 74.7 304 663 170
+SOFEPNet+VACKF+GMCS 64.2 76.2 75.4 286 689 169
+SOFEPNet+VACKF+GMCS+STMP 64.6 76.3 76.1 276 702 170
To address the limitations of classical Kalman filter in predicting nonlinear motion, we introduced several enhancements to the filtering process: a velocity-adaptive model was employed to capture the nonlinear dynamics of moving objects and was integrated into the CKF. As shown in Table 3, the inclusion of the velocity-adaptive model led to an increase in MOTA to 62.7 and IDF1 to 74.4, indicating the model’s ability to adapt to non-uniform motion and effectively mitigate estimation biases caused by conventional linear prediction assumptions. Building on this, the integration of the CKF resulted in the complete VACKF model, which further improved MOTA to 63.2 and IDF1 to 74.7, representing the highest scores in this group of experiments. These results demonstrate that CKF’s capability to model nonlinear state transitions complements the adaptive motion modeling, enhancing the filter’s ability to cope with complex and dynamic scenarios. Overall, the experimental results confirm that the design of VACKF significantly improves state estimation accuracy, reduces trajectory jitter and identity switches, and enhances the overall stability and robustness of the tracking system under challenging conditions.
Table 3: Performance test of VACKF module.
Speed Adaptive Model CKF MOTA $\uparrow$ IDF1 $\uparrow$
VACKF 62.5 74.1
62.7 74.4
63.2 74.7
MOT17: To further validate the generalization and robustness of the proposed modules, we integrated SocialTrack (excluding the specialized detector) into two mainstream multi-object tracking frameworks–ByteTrack [64] and BoT-SORT [1]–and conducted systematic ablation experiments on the MOT17 dataset while keeping the detector consistent (YOLOX-X). The experimental results are shown in Table 4 and Table 5.
Table 4: Performance improvement with different components added to ByteTrack.
@@
ByteTrack 68.0 76.6 79.3 75.1 84.1 159
ByteTrack+VACKF 68.1 76.7 79.7 75.4 84.4 160
ByteTrack+VACKF+GMCS 68.4 76.7 80.2 75.9 85.0 156
ByteTrack+VACKF+GMCS+STMP 68.3 76.7 80.4 76.1 85.3 139
Figure 7: Visual tracking results of the ByteTrack and the proposed SocialTrack in different scenarios in the MOT17 dataset.
In the ByteTrack framework, after adding VACKF, HOTA improved from 68.0 to 68.1, MOTA increased from 76.6 to 76.7, IDF1 rose from 79.3 to 79.7, and both IDR and IDP showed slight increases. This indicates that even under conditions of limited detection accuracy, VACKF can enhance identity preservation and overall stability in the tracking process through its nonlinear modeling capabilities and adaptive state estimation mechanism. After introducing the GMCS, the overall performance improved further, with IDF1 reaching 80.2, HOTA rising to 68.4, and IDSW decreasing from 159 to 156. By effectively modeling the motion consistency among individuals in the local neighborhood, this module demonstrated stronger identity continuity in scenarios involving occlusion and crossing motion. With the addition of the STMP, IDF1 further increased to 80.4, and IDSW significantly dropped to 139, demonstrating that this module can accurately predict states in situations with low observation quality or temporary target loss, thereby effectively mitigating identity switching and enhancing the system’s temporal robustness. Fig. 7 presents the tracking results between the ByteTrack and the proposed SocialTrack across various scenarios in the MOT17 dataset.
Table 5: Performance improvement with different components added to BoT-SORT.
@@
BoT-SORT 69.1 78.3 81.4 76.9 86.5 140
BoT-SORT+VACKF 69.1 78.4 81.4 76.9 86.5 140
BoT-SORT+VACKF+GMCS 69.3 78.4 81.7 76.9 87.0 123
BoT-SORT+VACKF+GMCS+STMP 69.4 78.4 82.2 77.4 87.6 122
Figure 8: Visual tracking results of the BoT-SORT and the proposed SocialTrack in different scenarios in the MOT17 dataset.
@@
In the BoT-SORT framework, a similar trend was observed, further confirming the good generalizability of the proposed method. Compared to the original BoT-SORT, VACKF did not lead to performance improvements, which can be attributed to the camera motion compensation already addressing the position prediction shortcomings. After introducing the GMCS, IDF1 improved from 81.4 to 81.7, IDP increased to 87.0, and IDSW significantly decreased to 123, indicating that this strategy can effectively integrate motion prior information into BoT-SORT’s feature modeling mechanism, improving association accuracy in weak observation scenarios. Finally, with the addition of the STMP, the IDF1 of BoT-SORT increased to 82.2, HOTA rose to 69.4, and IDSW further decreased to 122, showing a consistent performance improvement trend similar to ByteTrack. This module demonstrated robust and transferable performance enhancement across different frameworks, further proving the proposed method’s good cross-framework adaptability and potential for broad application. Fig. 8 presents the tracking results between the BoT-SORT and the proposed SocialTrack across various scenarios in the MOT17 dataset.
5 Conclusion
@@
This study proposes a novel multi-object tracking algorithm, SocialTrack, designed to address the challenges of tracking small targets in complex environments. Through a series of ablation and comparison experiments, we have demonstrated the effectiveness of the proposed improvements in enhancing tracking accuracy, robustness, and stability. Compared to current mainstream methods, SocialTrack leads in several metrics, particularly in core performance indicators such as MOTA (64.6%) and IDF1 (76.1%), achieving improvements of 9.3% and 4.8%, respectively, over the existing best methods, setting a new state-of-the-art. Additionally, SocialTrack shows superior performance in handling complex scenarios, particularly under frequent occlusions and dense targets, effectively reducing identity switching and target loss, demonstrating stronger robustness.
In conclusion, SocialTrack significantly improves the overall performance of multi-object tracking in complex environments by systematically integrating multiple innovative modules. Furthermore, all innovative modules possess strong plug-and-play capabilities and compatibility, allowing them to be incorporated into existing tracking algorithms to further enhance performance, demonstrating the great potential of this method for practical applications. Therefore, SocialTrack provides an important reference and insight for future multi-object tracking research and engineering applications.
Declarations
[…]
Acknowledgements This research received funding from the National Natural Science Foundation of China under Grant 62106193 and from the Fundamental Research Funds for the Central Universities under Grant G2024KY05105.
[…]
Author contributions Wenguang Tao: Data Curation, Conceptualization, Methodology, Software, Validation, Writing-Original Draft. Xiaotian Wang: Conceptualization, Software, Validation. Tian Yan & Guodong Li: Software, Validation, Visualization. Jie Yan: Supervision, Validation, Writing-Review & Editing. Kun Bai: Formal analysis, Supervision, Validation, Writing-Review & Editing. All authors reviewed the manuscript.
[…]
Data availability No datasets were generated or analysed during the current study.
[…]
Conflict of interest The authors declare no competing interests.